OpenStudy (anonymous):

solve for t: 0=t^2-2t-1

OpenStudy (blues):

Factor it.

jhonyy9 (jhonyy9):

t1,2=(2+/- sqrt8)/2 = 1+/-sqrt2

OpenStudy (anonymous):

what do you mean by t1,2?

jhonyy9 (jhonyy9):

those two roots

jhonyy9 (jhonyy9):

those two zeros of this function

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so do I put that there are no points?

jhonyy9 (jhonyy9):

check those hence (t -(1+sqrt2))(t-(1-sqrt2))

jhonyy9 (jhonyy9):

and you need getting this equation like result

OpenStudy (anonymous):

t^2-2t-1=(t-1)^2-2 now solve (t-1)^2=2

OpenStudy (anonymous):

that is how I was doing it.

jhonyy9 (jhonyy9):

yes but if you make those calcules what i have wrote you will get the same sure

OpenStudy (anonymous):

so there are many ways to do this problem????

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

yes i have doit those calcules by discriminant

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

am I back to being not sure about myself . Yes

OpenStudy (anonymous):

I have not done discriminants yet

jhonyy9 (jhonyy9):

so if you are not sure than check it by calcules

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

ok but you know how can being solved one quadratic equation by used discriminant ?

OpenStudy (anonymous):

somewhat yes

jhonyy9 (jhonyy9):

so we know that D=b2-4ac and t1,2 =(-b+/- sqrt(D))/2a right ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so now check it my firstly calcules and will can seeing the results

OpenStudy (anonymous):

my calc won't let me do it?

jhonyy9 (jhonyy9):

D=b2-4ac = 4 -4*1*(-1)=4+4=8 t1,2=(-b+/- sqrtD)/2a=(2+/-sqrt8)2=(2+/-2sqrt2)/2=2(1+/-sqrt2)/2= =1+/-sqrt2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yep

jhonyy9 (jhonyy9):

so than give medal

jhonyy9 (jhonyy9):

and good luck bye

OpenStudy (anonymous):

Thanks

jhonyy9 (jhonyy9):

yw